# Quantifying and Bridging the Fidelity Gap: A Decisive-Feature Approach to Comparing Synthetic and Real Imagery

### v1

@@ IV-B DFF Estimator
We instantiate DFF exclusively with a mask-and-infill counterfactual explainer [^1] Counterfactuals give intervention-style evidence and avoid well-known pitfalls of purely associative saliency, e.g., unrobustness/inconsistency of explanations [ghorbani2019interpretation, huang2023safari, zhao2021baylime]; we therefore choose CF–XAI. That said, conceptually any reliable XAI method can be applied in our DFF framework. (CF–XAI) [FongVedaldi2017Meaningful].
How CF-XAI works (intuition): Given an image $x$, the explainer asks: “What is the smallest change to $x$ that would flip the SUT’s output?” It finds a sparse mask $m^{\star}$ indicating which pixels must be modified (via infill from a learned prior) to change the decision. The resulting decisive map $\operatorname{H}(F(x))$ highlights these critical regions.
We compute $\operatorname{H}(F(x_{r}))$ for the real image and $\operatorname{H}(F(x_{s}))$ for the synthetic image using the same CF-XAI procedure applied independently. To ensure robustness against the stochastic nature of counterfactual generation, we define the decisive map as the expectation over $K_{\mathrm{cf}}$ random seeds $\zeta_{k}$:
$\operatorname{H}(F(x))\coloneqq\frac{1}{K_{\mathrm{cf}}}\sum_{k=1}^{K_{\mathrm{cf}}}m^{\star}(x;\zeta_{k}),$ (4)
where $m^{\star}(x;\zeta_{k})$ is the optimal mask generated for input $x$ using seed $\zeta_{k}$. We use mean squared error (MSE) as $\operatorname{D}$, computed on these averaged, pooled $16\times 16$ decisive maps.
@@ IV-C DFF-based Calibration Objective
where $x_{s,i}^{\ast}=G_{\Theta_{i}^{\ast}}(SD_{i})$, and $\mathcal{L}_{\text{recon}}$ and $\mathcal{L}_{\text{OV}}$ are standard reconstruction (for example $\ell_{1}$/LPIPS) and task losses (for example MSE). We define $\mathcal{L}_{\text{DFF}}(x_{s},x_{r})\coloneqq\operatorname{D}\bigl(\operatorname{H}(F(x_{s})),\operatorname{H}(F(x_{r}))\bigr)$. We treat $F$ as fixed and non-differentiable and backpropagate gradients only through $C_{\eta}$ into the generator knobs.
@@ IV-D Calibration Implementation
$C_{\eta}$ predicts generator/post-processing parameters per scenario. We optimize the random seed using Evolution Strategies (ES; population size 32, $\sigma{=}0.1$). For continuous parameters specifically ControlNet scales, guidance scale, and mild post-operations (contrast/brightness/blur) we use Stochastic Gradient Descent (SGD) with learning rates of $5\!\times\!10^{-3}$ and $1\!\times\!10^{-3}$ respectively. Gradients backpropagate through the calibrator and generator knobs only. We treat $F$ as fixed and non-differentiable. The DFF dissimilarity provides the signal for the optimization (SGD or ES) of these generator knobs.
@@
We optimize $C_{\eta}$ to minimize the combined loss in Eq. (5) using SGD/ES on the generator knobs; gradients do not pass through $F$. Decisive maps $\operatorname{H}(F(x_{r}))$ and $\operatorname{H}(F(x_{s}))$ are computed independently for each image in the pair using CF-XAI, then compared via $\operatorname{D}$ on pooled representations. Matched pairs or nearest neighbors are used, decisive sets are averaged across counterfactual (CF) seeds, and batching (i.e., computing explanations for $N$ images in parallel) amortizes the computational cost of the CF evaluations.
Algorithm 1 DFF-Guided Calibration Training
0: Real images $\{x_{r}\}$, segmentation maps $\{s\}$, frozen SUT $F$, ControlNet $G_{\Theta}$
0: Hyperparameters: steps T=22000T{=}22000, $\beta=0.3$, $\lambda_{\text{dff}}=0.08$, CF steps $K_{\mathrm{cf}}{=}80$
1:for $t=1$ to $T$ do
2:  Sample $(x_{r},s)$ from training set
3:  $x_{s}\leftarrow G_{\Theta}(s,\text{noise})$ {Generate synthetic image.}
4:  $\mathcal{L}_{\text{recon}}\leftarrow\text{MSE}(\text{predicted\_noise},\text{target\_noise})$
5:  $\mathcal{L}_{\text{OV}}\leftarrow\text{MSE}(F(x_{r}),F(x_{s}))$
6:  if $t\mod 3=0$ then
7:   $\operatorname{H}(F(x_{r}))\leftarrow\text{CF-XAI}(F,x_{r},K_{\mathrm{cf}})$ {Decisive map via Eq. 4}
8:   $\operatorname{H}(F(x_{s}))\leftarrow\text{CF-Optimize}(F,x_{s},K_{\mathrm{cf}})$ {Decisive map for synthetic}
9:   $\mathcal{L}_{\text{DFF}}\leftarrow\text{MSE}\bigl(\text{pool}(\operatorname{H}(F(x_{r}))),\text{pool}(\operatorname{H}(F(x_{s})))\bigr)$
10:  end if
11:  $\mathcal{L}_{\text{total}}\leftarrow\mathcal{L}_{\text{recon}}+\beta\cdot\mathcal{L}_{\text{OV}}+\lambda_{\text{dff}}\cdot\mathcal{L}_{\text{DFF}}$
12:  Update $\Theta$ via backpropagation on $\mathcal{L}_{\text{total}}$
13:end for
14:return Calibrated generator $G_{\Theta^{*}}$
@@ V-A Experimental Setup
Data and pairing. We evaluate on 2,126 matched KITTI–VirtualKITTI2 pairs across the five VKITTI2 scenes (Scene01, 02, 06, 18, 20), split 80/20 by scene into 1,701 calibration and 425 held-out items.
VirtualKITTI2 provides synthetic “clone” sequences that follow the same camera trajectory as the corresponding KITTI raw sequences. We form KITTI–VKITTI2 pairs by matching the scene ID, rendering variant (VirtualKITTI2’s “clone” condition replicates KITTI lighting and weather), camera (Camera_0), and frame index.
SUTs. We use three fixed SUT heads: (i) a PilotNet-style steering angle prediction CNN; (ii) YOLOP drivable-area; (iii) YOLOP lane-line. Feature layers are conv5 (steering) and neck pre-decoder (YOLOP), selected for semantic relevance. All SUT weights are frozen (i.e., the SUTs are not trained or fine-tuned; we evaluate their fixed, pre-trained behavior).
Metrics. Details are given in Section V-B.
Calibration. We optimize generator knobs with a combined loss on IV, OV, and DFF, treating the SUT as fixed.
Full details. Optimizers, CF-XAI hyperparameters, and per-layer shapes are documented in the anonymous repository ((link hidden for review)).
Fidelity thresholds were selected based on domain standards and empirical distributions: $\varepsilon_{\mathrm{out}}$ corresponds to a high-fidelity cutoff of 0.7 similarity (e.g., $\text{IoU}\geq 0.7$), while $\varepsilon_{\mathrm{dff}}$ is set to 0.3 based on the observed loss distribution of the explainer. Appendix A summarizes the notation used in the paper.
@@ V-B Metrics and thresholds
Recall from Eq. (1) that IV fidelity is defined by a distance metric $\operatorname{D_{\mathrm{in}}}$. We instantiate $\operatorname{D_{\mathrm{in}}}$ using LPIPS [Zhang2018LPIPS] (lower is better). To facilitate visual comparison in plots where "higher is better," we define a normalized IV Score as $1-\text{LPIPS}_{\text{norm}}\in[0,1]$, where LPIPS is clamped to $[0,1]$. This score is used solely for visualization; the rigorous fidelity check uses the raw LPIPS distance against $\varepsilon_{\text{in}}$.
@@
Full backbone choices, normalization, non-inferiority (NI) margins, and inference details are documented in the anonymous repository ((link hidden for review)).
@@ V-C Results and Analysis
Figure 2: Decisive-Feature Fidelity across SUTs on the same dataset. (a) Distribution of DFF distances $D$ for Steering, YOLOP–DA, and YOLOP–$L$; vertical dashed lines mark each SUT’s $\varepsilon_{95}$ threshold. (b) Empirical CDFs $P(D\leq\varepsilon)$; the horizontal dashed line indicates 95% coverage, and markers show where each CDF crosses this level. The three curves differ substantially, confirming that DFF is SUT-specific: heads processing the same frames exhibit different decisive-feature mismatch profiles. While OV losses are minimal (see Table III), DFF retains substantial spread, indicating mechanism differences not captured by IV/OV.
@@
[…]
@@
For Steer, optimizing only OV loss (OVF-Steer) improves LPIPS (0.610 $\to$ 0.473) and DFF (0.361 $\to$ 0.278), but dramatically worsens OV loss (0.0006 $\to$ 0.3908). This occurs because the OVF optimization can overfit to perceptual similarity at the expense of task-relevant output alignment. The DFF-Steer variant achieves the best DFF (0.229) while preserving OV loss near baseline levels (0.0028), demonstrating that DFF-guided calibration successfully aligns decisive features without sacrificing output fidelity.
For DA, OVF improves OV loss (0.0172 $\to$ 0.0044) but worsens DFF (0.541 $\to$ 0.630), illustrating the core problem: output-value parity without mechanism parity. The SUT produces similar outputs but relies on different decisive features. In contrast, DFF-DA achieves improved OV loss (0.0172 $\to$ 0.0044) with a smaller DFF increase than OVF-DA (0.541 $\to$ 0.577 vs. 0.541 $\to$ 0.630), showing that aligning decisive features can mitigate mechanism divergence while benefiting output fidelity.
For LL, both calibration variants improve over baseline. OVF-LL achieves the best OV loss (0.0019 $\to$ 0.0011) and also improves DFF (0.712 $\to$ 0.630). The DFF-LL variant achieves competitive DFF (0.635) with similar OV loss (0.0013). For this SUT, OVF-only calibration yields slightly better DFF than the DFF-guided variant; this pattern persists in the aggregate (Table V), though both variants improve over baseline. This may reflect that for lane-line detection, optimizing output agreement incidentally aligns the decisive features the SUT uses.
These patterns support O1: OV loss can improve while DFF worsens (DA, OVF variant), demonstrating output-value parity without mechanism parity. DFF-guided calibration achieves the lowest DFF for Steering and DA while maintaining competitive OV loss; for LL, OVF-only calibration happens to yield marginally better DFF, though both variants improve substantially over baseline. The DA case most clearly illustrates the failure mode DFF is designed to reveal: OVF-DA reduces OV loss by 74% yet increases DFF by 16%, meaning the SUT produces more similar outputs but relies on more divergent decisive features.
The held-out distribution and CDF of DFF distances show clear SUT-specific variation while OV remains tightly clustered (Fig. 2; lower is better for DFF). The weak coupling is quantified by Spearman correlations in Table IV, confirming DFF carries information not captured by IV/OV.
@@
Figure 3: Qualitative comparison of calibration methods on a representative KITTI–Virtual KITTI 2 pair. (a) Real KITTI input. (b) Uncalibrated baseline synthetic. (c)–(e) OVF-calibrated outputs for Steering, Drivable Area, and Lane Lines respectively. (f)–(h) DFF-calibrated outputs for the same three SUTs. Each synthetic panel displays its IV score (labeled “IV”, computed as $1-\text{LPIPS}$), OV score, and DFF distance (labeled “D”). Cyan boxes (Region A) highlight road-surface texture; magenta boxes (Region B) highlight sky and horizon. DFF calibration produces more realistic road texture and atmospheric rendering in the decisive regions, whereas OVF calibration optimises output similarity without necessarily improving perceptual realism in SUT-critical areas.
@@
Figure 3 summarizes the contrast between optimizing only for outputs and optimizing for decisive features. The OVF variants (rows 3–5) reproduce the SUT outputs on the matched scene, yet their visual evidence differs in ways that can shift the SUT’s internal reasoning. The DFF-calibrated variants (rows 6–8) keep outputs comparable while making the cues that drive the decision, such as road boundaries, lane markings, and vehicle contours, more consistent with the real frame. This aligns with our O1 claim that output similarity does not guarantee mechanism parity, and it motivates using DFF alongside IV and OV in assessment.
@@
Table V summarizes the core effect sizes on the held-out split: DFF-guided calibration reduces DFF (negative $\Delta\text{DFF}$, lower is better) while leaving OV loss non-inferior and slightly improved on average (positive $\Delta\text{OV}$). Non-inferiority holds under the pre-declared margins (steering $\delta{=}-0.005$, YOLOP $\delta{=}-0.010$), corresponding to roughly 0.5% and 1% relative degradation in OV score. These thresholds were chosen a priori from pilot experiments: we identified the smallest OV drop that visibly affected downstream behavior (e.g., lane-keeping wobble), then set $\delta$ conservatively below that level.
@@ VI Threats to Validity
Qualitative example selection. The frame shown in Fig. 3 contains faint or partially occluded lane markings, which may seem suboptimal for evaluating YOLOP-LL. We deliberately retained this frame because (i) it is representative of the dataset’s natural difficulty distribution, and (ii) it illustrates that DFF calibration improves decisive-feature alignment even when ground-truth lane cues are weak—a practically relevant regime where spurious evidence is most likely to mislead the SUT.
Decisive-set identification. We use a mask-and-infill CF–XAI estimator where each image (real or synthetic) serves as its own reference for counterfactual perturbation. The optimizer enforces sparsity and plausibility constraints, and we average over multiple CF random seeds to reduce variance. We compute $\operatorname{D}$ on $16{\times}16$ pooled decisive maps. We include stability checks (e.g., model-randomization sanity tests) to guard against explanation artifacts.

### v3

@@ IV-B DFF Estimator
We instantiate DFF exclusively with a mask-and-infill counterfactual explainer [^1] Counterfactuals give intervention-style evidence and avoid well-known pitfalls of purely associative saliency, e.g., unrobustness/inconsistency of explanations [9, 11, 29]; we therefore choose CF–XAI. That said, conceptually any reliable XAI method can be applied in our DFF framework. (CF–XAI) [7]. Intuitively CF-XAI works as follows: Given an image $x$, the explainer asks: “What is the smallest change to $x$ that would affect the SUT’s output?” It finds a sparse mask $m^{\star}$ indicating which pixels must be modified (via infill from a learned prior) to change the decision. The resulting decisive map $\operatorname{H}(F(x))$ highlights these critical regions.
We compute $\operatorname{H}(F(x_{r}))$ and $\operatorname{H}(F(x_{s}))$ independently with the same CF-XAI procedure and define the decisive map by multi-seed averaging:
$\operatorname{H}(F(x))\coloneqq\frac{1}{K_{\mathrm{cf}}}\sum_{k=1}^{K_{\mathrm{cf}}}m^{\star}(x;\zeta_{k}).$ (5)
We set $\operatorname{D}$ to MSE on pooled $16{\times}16$ averaged maps.
@@ IV-C DFF-based Calibration Objective
where $x_{s,i}^{\ast}=G_{\Theta_{i}^{\ast}}(SD_{i})$, and $\mathcal{L}_{\text{recon}}$ and $\mathcal{L}_{\text{OV}}$ are standard reconstruction (for example $\ell_{1}$/LPIPS) and task losses (for example MSE). We define $\mathcal{L}_{\text{DFF}}(x_{s},x_{r})\coloneqq\operatorname{D}\bigl(\operatorname{H}(F(x_{s})),\operatorname{H}(F(x_{r}))\bigr)$. We treat $F$ as fixed and non-differentiable and backpropagate gradients only through $C_{\eta}$ into the generator parameters.
@@ IV-D Calibration Implementation
$C_{\eta}$ predicts generator/post-processing parameters per scenario. We optimize the random seed using Evolution Strategies (ES; population size 32, $\sigma{=}0.1$). For continuous parameters specifically ControlNet scales, guidance scale, and mild post-operations (contrast/brightness/blur) we use Stochastic Gradient Descent (SGD) with learning rates of $5\!\times\!10^{-3}$ and $1\!\times\!10^{-3}$ respectively. Gradients backpropagate through the calibrator and generator parameters only. We treat $F$ as fixed and non-differentiable. The DFF dissimilarity provides the signal for the optimization (SGD or ES) of these generator parameters.
@@
We optimize $C_{\eta}$ to minimize the combined loss in  (6) using SGD/ES on the generator parameters; gradients do not pass through $F$. Decisive maps $\operatorname{H}(F(x_{r}))$ and $\operatorname{H}(F(x_{s}))$ are computed independently for each image in the pair using CF-XAI, then compared via $\operatorname{D}$ on pooled representations. Matched pairs or nearest neighbors are used, decisive sets are averaged across counterfactual (CF) seeds, and batching (i.e., computing explanations for $N$ images in parallel) amortizes the computational cost of the CF evaluations.
Algorithm 1 Training the calibrator $C_{\eta}$ (DFF-guided)
0:  Paired data $\{(x_{r},\,x_{s}^{(\mathrm{init})},\,SD)\}$, generator $G_{\Theta}$, frozen SUT $F$, explainer $\operatorname{H}$
1:  Initialize calibrator $C_{\eta}$
2:  repeat
3:   Sample a mini-batch $\mathcal{B}$
4:   for each $(x_{r},x_{s}^{(\mathrm{init})},SD)\in\mathcal{B}$ do
5:    $\Theta^{*}\leftarrow C_{\eta}(x_{s}^{(\mathrm{init})},SD)$
6:    $x_{s}^{*}\leftarrow G_{\Theta^{*}}(SD)$
7:    $L\leftarrow L_{\mathrm{IV}}(x_{s}^{*},x_{r})+L_{\mathrm{OV}}(F(x_{s}^{*}),F(x_{r}))+\operatorname{D}(\operatorname{H}(F(x_{s}^{*})),\operatorname{H}(F(x_{r})))$
8:   end for
9:   Update $\eta$ to minimize the batch loss (keeping $F$ fixed)
10:  until convergence
11:  Output: trained parameters $\eta^{*}$
@@ V-A Experimental Setup
Data and pairing. We evaluate on 2,126 matched KITTI–VirtualKITTI2 pairs across the five VKITTI2 scenes (Scene01, 02, 06, 18, 20), split 80/20 by scene into 1,701 calibration and 425 held-out items. VirtualKITTI2 provides synthetic “clone” sequences that follow the same camera trajectory as the corresponding KITTI raw sequences. We form KITTI–VKITTI2 pairs by matching the scene ID, rendering variant (VirtualKITTI2’s “clone” condition replicates KITTI lighting and weather), camera (Camera_0), and frame index. SUTs. We use three fixed SUT heads: (i) a PilotNet-style steering angle prediction CNN; (ii) YOLOP drivable-area (DA); (iii) YOLOP lane-line (LL). Feature layers are conv5 (steering) and neck pre-decoder (YOLOP), selected for semantic relevance. All SUT weights are frozen (i.e., the SUTs are not trained or fine-tuned; we evaluate their fixed, pre-trained behavior). Metrics. Details are given in Section V-B. Calibration. We optimize generator parameters with a combined loss on IV, OV, and DFF, treating the SUT as fixed. Full details. Optimizers, CF-XAI hyperparameters, and per-layer shapes are documented here: (DFF-repo). Fidelity thresholds follow domain standards and calibration distributions: $\varepsilon_{\mathrm{out}}$ uses a 0.7 similarity cutoff (e.g., $\mathrm{IoU}\!\geq\!0.7$).
@@ V-B Metrics and thresholds
Recall from Eq. (1) that IV fidelity is defined by a distance metric $\operatorname{D_{\mathrm{in}}}$. We instantiate $\operatorname{D_{\mathrm{in}}}$ using LPIPS [27] (lower is better). To facilitate visual comparison in plots where "higher is better," we define a normalized IV Score as $1-\text{LPIPS}_{\text{norm}}\in[0,1]$, where LPIPS is clamped to $[0,1]$. This score is used solely for visualization; the rigorous fidelity check uses the raw LPIPS distance against $\varepsilon_{\text{in}}$.
@@
Full backbone choices, normalization, non-inferiority (NI) margins, and inference details are documented in the repository (DFF-repo).
@@ V-C Results and Analysis
Figure 2: DFF across SUTs. (a) DFF distance $D$ distributions; dashed lines: $\varepsilon_{95}$. (b) CDFs $P(D\leq\varepsilon)$; dashed line: 95%. Curves differ substantially, confirming SUT-specific mechanism mismatch even when OV losses are small (Table III).
@@
TABLE III: Pilot test on one matched pair (Scene20, frame 750). Lower is better for LPIPS and DFF. For OV, we report raw task loss: MSE for steering and $(1{-}\text{IoU})$ for YOLOP heads; lower loss indicates better output agreement. OVF-Steer’s high OV loss reflects overfitting to perceptual similarity at the expense of task output.
@@
For Steer, optimizing only OV loss (OVF-Steer) improves IV loss (0.610 $\to$ 0.473) and DFF (0.361 $\to$ 0.278), but dramatically worsens OV loss (0.0006 $\to$ 0.3908). This occurs because the OVF optimization can overfit to perceptual similarity at the expense of task-relevant output alignment. The DFF-Steer variant achieves the best DFF (0.229) while preserving OV loss near baseline levels (0.0028), demonstrating that DFF-guided calibration successfully aligns decisive features without sacrificing output fidelity. For DA, OVF improves OV loss (0.0172 $\to$ 0.0044) but worsens DFF (0.541 $\to$ 0.630), illustrating the core problem: output-value parity without mechanism parity. The SUT produces similar outputs but relies on different decisive features. In contrast, DFF-DA achieves improved OV loss (0.0172 $\to$ 0.0044) with a smaller DFF increase than OVF-DA (0.541 $\to$ 0.577 vs. 0.541 $\to$ 0.630), showing that aligning decisive features can mitigate mechanism divergence while benefiting output fidelity. For LL, both calibration variants improve over baseline. OVF-LL achieves the best OV loss (0.0019 $\to$ 0.0011) and also improves DFF (0.712 $\to$ 0.630). The DFF-LL variant achieves competitive DFF (0.635) with similar OV loss (0.0013). For this SUT, OVF-only calibration yields slightly better DFF than the DFF-guided variant; this pattern persists in the aggregate (Table V), though both variants improve over baseline. This may reflect that for lane-line detection, optimizing output agreement incidentally aligns the decisive features the SUT uses. These patterns support O1: OV loss can improve while DFF worsens (DA, OVF variant), demonstrating output-value parity without mechanism parity. DFF-guided calibration achieves the lowest DFF for Steering and DA while maintaining competitive OV loss; for LL, OVF-only calibration happens to yield marginally better DFF, though both variants improve substantially over baseline. The DA case most clearly illustrates the failure mode DFF is designed to reveal: OVF-DA reduces OV loss by 74% yet increases DFF by 16%, meaning the SUT produces more similar outputs but relies on more divergent decisive features. The held-out distribution and CDF of DFF distances show clear SUT-specific variation while OV remains tightly clustered (Fig. 2; lower is better for DFF). The weak coupling is quantified by Spearman correlations in Table IV, confirming DFF carries information not captured by IV/OV.
@@
Figure 3: Qualitative comparison of calibration methods on a representative KITTI–Virtual KITTI 2 pair. (a) Real KITTI input. (b) Uncalibrated baseline synthetic. (c)–(e) OVF-calibrated outputs for Steering, Drivable Area, and Lane Lines respectively. (f)–(h) DFF-calibrated outputs for the same three SUTs. Each synthetic panel displays its IV score (labeled “IV”, computed as $1-\text{LPIPS}$), OV score, and DFF distance (labeled “D”). Cyan boxes (Region A) highlight road-surface texture; magenta boxes (Region B) highlight sky and horizon.
@@
Figure 3 illustrates the contrast between optimizing only for outputs and optimizing for decisive features. The OVF variants (second row) reproduce the SUT outputs on the matched scene, yet their visual evidence differs in ways that can shift the SUT’s internal reasoning. DFF calibration produces more realistic road texture and atmospheric rendering in the decisive regions, whereas OVF calibration optimises output similarity without necessarily improving perceptual realism in SUT-critical areas. The DFF-calibrated variants (third row) keep outputs comparable while making the cues that drive the decision, such as road boundaries, lane markings, and vehicle contours, more consistent with the real frame.
The frame in Fig. 3 contains faint or partially occluded lane markings, which is challenging for YOLOP–LL. We retained this case because it is representative of the dataset’s natural difficulty distribution, and it highlights that DFF-guided calibration can improve decisive-feature alignment even when lane cues are weak, a regime where spurious evidence is more likely to mislead the SUT.
@@
Table V summarizes the core effect sizes on the held-out split: DFF-guided calibration reduces DFF (negative $\Delta\text{DFF}$, lower is better) while OV loss is slightly improved on average (positive $\Delta\text{OV}$). Non-inferiority holds under the pre-declared margins (steering $\delta{=}-0.005$, YOLOP $\delta{=}-0.010$), corresponding to roughly 0.5% and 1% relative degradation in OV score. These thresholds were chosen a priori from pilot experiments using a conservative procedure: we swept calibration settings to induce controlled OV-score drops, identified the smallest drop that produced a consistent and visible downstream degradation in a simple lane-keeping check, and then set $\delta$ strictly below that value.
@@ VI Threats to DFF Validity
Qualitative illustration limitations. Fig. 3 is provided as an illustrative example to support intuition, but any single qualitative case may be unrepresentative. Therefore, our conclusions are primarily supported by the aggregate quantitative results on the held-out split; the qualitative figure is not used as standalone evidence.
Explanation noise. CF–XAI is stochastic, which adds noise to DFF and to calibration feedback. We reduce variance via multi-seed averaging (Eq. 5), pooling, and sanity checks, and report sensitivity to $K_{\mathrm{cf}}$.
Decisive-set identification. We use mask-and-infill CF–XAI with sparsity/plausibility constraints, average over CF seeds, and compare $16{\times}16$ pooled maps; stability checks (e.g., model-randomization) help guard against artifacts.